OpenStudy (anonymous):

Solve using substitution method. 6x-3y=18, 6x+3y=-12

OpenStudy (anonymous):

This is what I have done so far not sure if it is right. 6x-3y=18 6x+3y=-12 12x=-12 x=-1 6(-1) -3Y=18 -6-3Y=18 +6 +6 -3y=24 /-3 /-3 y=-8

OpenStudy (anonymous):

Using the second equation: -12 - 6x = 3y -4 - 2x = y So 6x - 3(-4 - 2x) = 18 6x + 12 + 6x = 18 12x = 6 x = 1/2 Since -4 - 2x = y - 4 - 2(1/2) = y -4 - 1 = y y = -5 Check your solution: 6x - 3y = 18 6(1/2) - 3(-5) = 18 3 + 15 = 18 18 = 18 Yup, it checks out. So x = 1/2 and y = -5!